Mars Society Flashline Arctic Research Station Mission Support
Home About Team Mission Logs FMARS Home RMMS Home

FMARS Mission Support Summary-Sunday, July 13th

5 pm CDT July 13th, 2001 Status Report from FMARS field team, second rotation

Our status is thus;

EVA planned for today to weather station was prevented by rain and high water in the Lowell Canal.

We spent the day in the hab writing reports and preparing EVA gear.

The power shortage is bothersome, but livable. We cannot use the electric heaters, so temperature in the hab varies between 50 and 60 F.

Water usage between 8 am July 12 and 9 am July 13, was only 14 gallons, which compares to 48 gallons for 6 people estimated by NASA. However we then allowed two crew members to take brief "Navy" showers, and an additional 13 gallons was used.

Everyone is in good spirits and good health. Weather seems to be improving. Assuming the trend continues, we plan a four person motorized EVA tomorrow to Trinity Lake and the Breccia Hills for exobiology samples. EVA will be led by Charles, and include Katy, Vladimir, and Bill.

The main technical problem continues to be getting Steve's hi-speed data link radio system on line. The analyzer arrived at camp today, but has not yet been transported here due to high water in the Lowell Canal through most of the day. The water level is now lower, and we hope to have the analyzer soon. If so Steve will be able to say in 24-48 hours if he will be able to get his dish working soon, or if another week will be required while a new amplifier is brought in.

We made suit #8, designed for tall people, operational today. We have also broken out the hi-gain suit radio antennas, and will implement them on our next EVA.

Robert Zubrin

STATION Vladimir Pletser (report has been abridged by Mission Support)

A three member EVA crew (Robert Zubrin, Katy Quinn, Vladimir Pletser) conducted a seismic experiment during a four hour EVA in poor weather conditions (rain and wind); see Figure 1. A line of 24 seismic sensors (geophones) were deployed in the Haynes Ridge plain, in front of the Flashline MARS Habitat. The line (geophone flute) was laid in a South-South-East direction, approximately parallel to the Haughton crater rim. This direction was chosen as being perpendicular to the direction of the flute laid in a previous dry run trial. The geophones were spaced every 4 m. Three tests were conducted with a trigger geophone triggered by quakes created by sledge hammer shots. The trigger geophone and the hammer shots were respectively located at the middle of the flute (Test # 1) and at both ends of the flute (Tests # 2 and 3). Locations were measured using a hand-held GPS receiver. The geophones were tested individually and automatically prior to conducting the first Test and were found in functioning conditions. A first glance analysis of data back in the Hab showed the following:

  1. For the three test configurations, the average underground velocity of the signal was approximately 2600 m/s.
  2. Sounding extended to vertical depths in excess of 550 m.

In conclusion, no water is detected under the ground of Haynes Ridge, Average velocities of wave transmission for water are catalogued between 1450 and 1500 m/s for liquid water and between 3300 and 3800 m/s for ice. The average velocity deduced from test results is consistent with Calcium Carbonate and Dolomite (catalogued range: 12000 to 7000 m/s), which is commonly found in this area. More detailed analysis will be conducted after the campaign to further characterize the underground structure.

Chief Flight Engineer Steve Braham's Report: (Report has been abridged by Mission Support)

SFU/CSA/SETI Institute/NASA High Speed Field Comms The high-speed, PlanetNet 1, system continues to function perfectly. The Hab-based comms node has been moved downstairs, and a conventional IEEE 802.11 wireless network deployed on the upper deck, allowing the crew to have wireless network access from any location on that deck, including in their staterooms.

Power Systems The 1800W generator is functioning well, powering on-board comms and computing systems. UPS will be brought up soon to support critical components. The 5000W system continues to trip out, and brown out, which knocks out less fragile systems.

Water Systems Water pumping was not a problem today. However, the commander's experimental protocol for the incinerating toilet (no use of incineration function during sleeping hours) resulted in overflow of the storage container. The toilet has been cleaned out, and the CFE sterilized the area with large amounts of bleach.

ATVs We presently have 1 ATV.

12:30 PM CDT July 13th

Mission Support: We had a meeting late yesterday afternoon to discuss plans for the remaining 5 days of the shift.

Forecast is for bad weather on the 13th, followed by mixed, but somewhat bettwe weather in the following days.

So we have planned a short 2 person EVA led by Bill Clancey to administer a weather station on the 13th. We may have to cancel this if the weather does not improve.

On the 14th, assuming good weather, we will conduct a 4 person motorized EVA to Trinity Lake and the Breccia Hills. EVA will be led by Charles Cockell, and include Katy, Vladimir, and Bill.

On the 15th, we will conduct a motorized recon EVA to find a scientifically interesting site to deploy the geophone flute. EVA will be led by Charles, and include RZ, Vladimir, and Katy

On the 16th, we will conduct a motorized EVA to deploy the geophone at the slected remore site. EVA will be led by Vladimir, and include RZ, Katy, and Bill Clancey.

The 17th is margin in case the above schedule is pushed to the right.

We anticpate crew rotation at 9 pm local time on the 17th.

Robert Zubrin

Requests concerning this web site should be directed to the Webmaster.
Mars Illustration Copyright 1999 by Mike Koonce/Mars Society
Copyright © 2001 The Mars Society. All rights reserved.
Mars Society Flashline Arctic Research Station Mission Support